20L of So2 diffuse through a porous partition in 60seconds.what volume of oxygen will diffuse under similar condition in 30seconds?

Grade:12th pass

2 Answers

Shirin Poddar
6 years ago
r1/r2 = M2^1/2 / M1^1/2 (Grahams law of diffusion) Or, (20/60) / ( v/30) = 32^1/2 / 64^1/2 Or, 10/v = 1/root 2 Or, V=10 × root 2 Or, V= 14.14 L
jai
6 years ago
The answer is 14.14 Litres.Given: Volume of SO2=20 liter, Time taken for diffusion by SO2= 60 sec, Time taken for diffusion by O2= 30 sec.To find: Volume of O2Solution:Let the volume of O2 be VRate of diffusion of SO2=20/60=1/3Rate of diffusion of O2=V/30We know, r1/r2=root of m2/root of m1 (where m1 and m2 are the molecular masses)1/3/V/30=square root of 32/square root of 641/3* root 2=V/3010*root 2=V10*1.414=V14.14 litres=Vi.e Volume of O2 is 14.14 litres